1 / 5

October 2011 Experimenters Meeting: Accelerator Operation Summary and Status Report

This report summarizes the accelerator operations from October 3 to October 10, 2011, during Calendar Week #40. It includes weekly integrated intensity metrics for NuMI (7.59E18 protons), BNB (5.25E18 protons), and MTest (8.08E13 protons), along with recorded downtime of 69.52 hours. Key completed work and plans encompass proton source repairs, LRF1&2 tank vacuum leak repairs, BRF8 cavity removal, and ongoing safety system tests. Updates on Pbar production, and issues regarding cryo-compressors are also addressed.

trish
Télécharger la présentation

October 2011 Experimenters Meeting: Accelerator Operation Summary and Status Report

An Image/Link below is provided (as is) to download presentation Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author. Content is provided to you AS IS for your information and personal use only. Download presentation by click this link. While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server. During download, if you can't get a presentation, the file might be deleted by the publisher.

E N D

Presentation Transcript


  1. All Experimenters’ Meeting October 10, 2011

  2. Accelerator Operation Summary From : October 03, 2011 (0000 hours) To: October 10, 2011 (0000 hours) Calendar Week # 40 NuMI Weekly Integrated Intensity 7.59E18 protons BNB Weekly Integrated Intensity 5.25e18 protons MTest Weekly Integrated Intensity 8.08e13 protons Recorded Downtime 69.52 hours

  3. Status /Work Completed/Plans • Proton Source • LRF1&2 tank vacuum leak repaired. • LRF5 PA replacement • Removal of BRF8 cavity. • PA replacement on 3 BRF stations • Booster Safety system tests • Main Injector • ANU pre-shutdown jobs. • Safety system tests . • $23 ramp length changed on Thursday • 2.06 Second Rep rate in MI • Pbar • Terminated final Pbars on Friday Oct. 7 @ 1117. • Valved out EQ923 due to LCW leak. • NuMI • Target moved to LE10 • MiniBooNE • MI12 AC unit will be installed when it arrives. • MTA • Investigating breaker trips affecting cryo-compressors

More Related